OpenStudy (anonymous):

Each side of a square is increasing at a rate of 6cm/s. At what rate is the area of the square increasing when the area of the square is 16cm^(2) Please show the steps so I can see how to answer questions such as these

OpenStudy (anonymous):

The answer is 48cm^(2) I'm not even getting results close to this

OpenStudy (jamesj):

Let x = length of side of the square. Then area of the square is A(x) = x^2 You are also told that dx/dt = 6 cm/s Now, you want to find dA/dt when A = 16, i.e., when x = 4. By the chain rule \[ \frac{dA}{dt} = \frac{dA}{dx}\frac{dx}{dt} \] Try that.

OpenStudy (anonymous):

I'm still lost I understand what dA/dt is but I just dont get how to solve it dA/dt A' = 2xx' 16/2(6) = x' x' = 16/12 which is incorrect what am I doing wrong.

OpenStudy (anonymous):

yeah I'm still completely lost

OpenStudy (anonymous):

where are you getting x = 4 from

OpenStudy (anonymous):

Ok I get it now

OpenStudy (jamesj):

so tell me

OpenStudy (anonymous):

Since dx/dt = 6cm/s and dA/dx = 16cm/s dA/dx dx/dt = dA/dt dA/dt A' = 2xx' A' = 2x6cm/s So we need to solve for x because we are trying to find dA so 16 = x^(2) (16)^(1/2) = 4 thus A' = 2(4)(6) A' = 48

OpenStudy (anonymous):

Thanks for the help, I think I finally understand d/dx now

OpenStudy (jamesj):

No, no. Wait.

OpenStudy (anonymous):

don't worry I'm not going anywhere did I do something wrong?

OpenStudy (jamesj):

By the chain rule \[ \frac{dA}{dt} = \frac{dA}{dx} \frac{dx}{dt} \] Look at the function for A, it is A(x) = x^2. Hence dA/dx = 2x. Further, you are told that dx/dt = 6. Therefore using the equation above \[ \frac{dA}{dt} = \frac{dA}{dx} \frac{dx}{dt} = 2x . 6 = 12x \]

OpenStudy (jamesj):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

When the area of the square is 16, then x = 4. Hence instantaneously, at the moment, \[ \frac{dA}{dt} = 12x = 12(4) = 48 \]